1. A worker requires 8 tanks of crude oil and petrol to get a work done. He gains 5000 

Ghana cedis per tank of crude oil used and 3000 Ghana cedis per petrol used for any 

work done. To finish the work, he further requires 10 tanks by combining 2 tanks of 

crude oil and a tank of petrol. What is the maximum profit he can make


1
Expert's answer
2021-04-14T14:18:50-0400

Let,

x = number of tanks of crude oil

y = number of tanks of petrol

We have to maximize "Z = 5000x + 3000y" ---- (1)

Subject to the conditions,

"2x + y = 10"

"x+y = 8"

After solving these equations we get "x=2" and "y = 6"

Substituting the values of x and y in equation (1)

We get,

"Z = 5000(2)+3000(6)"

"Z = 280000"